You have a portfolio that has 125 shares of Stock A that sell for $54 per share and 112 shares of Stock B that sell for $93 per share. What is the portfolio weight for Stock A? ENTER YOUR ANSWER AS A PERCENTAGE WITH ONE DECIMAL PLACE (e.g., 12.1) AND NOT AS A DECIMAL (e.g., 0.121). ROUND TO THE NEAREST TENTH OF A PERCENT.
Answer #1

Total value of:

A=(125*54)=6750

B=(112*93)=10416

Total value=(6750+10416)=17166

Hence weight of A=6750/17166

=39.3%(Approx).